# Orthogonal Howe duality and dynamical (split) symmetric pairs

### v2

@@ 8.1. Background on connections and differential operators
A flat connection is a conection $\nabla$ such that $[\nabla_{\xi},\nabla_{\eta}]=\nabla_{[\xi,\eta]}$, for all vector fields $\xi,\eta$.
In coordinates, $\nabla$ is a flat connection if and only if $[\nabla_{\partial/\partial z_{i}},\nabla_{\partial/\partial z_{j}}]=0$ for all $1\leq i,j\leq n$.
@@ 8.2. Boundary Casimir connection
In this subsection, we are in the setup of Section 5. For instance, we consider an arbitrary simple Lie algebra $\mathfrak{g}$ with a Cartan involution $\theta$.
@@
One can easily see that Theorem 8.7 follows from the following proposition.
@@
Let $N_{\alpha,\gamma}\in\mathbb{C}$ be the numbers such that $[E_{\alpha},E_{\gamma}]=N_{\alpha,\gamma}E_{\alpha+\gamma}$. Then a summand of the expression above is equal to
@@
$[T(\lambda,\mu),T(\nu,\mu)]=\sum_{\alpha,\beta\in R_{+}}\frac{\langle\alpha,\lambda\rangle\langle\beta,\nu\rangle}{\langle\alpha,\mu\rangle\langle\beta,\mu\rangle}[B^{2}_{\alpha},B^{2}_{\beta}]=\sum_{\begin{subarray}{c}\alpha,\beta\in R_{+}\\ \alpha\neq\beta\end{subarray}}\frac{\langle\alpha,\lambda\rangle\langle\beta,\nu\rangle}{\langle\alpha,\mu\rangle\langle\beta,\mu\rangle}[B^{2}_{\alpha},B^{2}_{\beta}].$
@@
The residue [^6] Recall for example that if $f$ is a rational function on $\mathbb{C}^{n}$, the residue along $z_{1}=0$ is computed as follows. Write $f=z_{1}^{-k}f_{-k}+z_{1}^{-k+1}f_{-k+1}+\dots$, where each $f_{i}$ is a rational function of $z_{2},\dots,z_{n}$, then $\mathrm{res}_{z_{1}=0}f=f_{-1}$. In particular, if $f$ has only a simple pole at $z_{1}=0$, i.e. $f=z_{1}^{-1}f_{-1}+f_{0}+z_{1}f_{1}+\dots$, then $\mathrm{res}_{z_{1}=0}f=z_{1}|_{z_{1}=0}$. at $\langle\alpha,\mu\rangle=0$ is given by
@@
by 8.9. Since the sum in equation (8.3) can presented as the sum of such terms, the residue is zero. We deduce that the function is regular.
It follows from Proposition 3.5 and 3.15 that
@@
In particular, by Proposition 8.4, we see that the boundary Casimir connection descends to the quotient $\mathfrak{h}^{\mathrm{reg}}/W$. Recall that the monodromy of the usual Casimir connection is given by the quantum Weyl group action [ToledanoLaredoQuantumWeyl]. Analogously, we propose the following conjecture.
@@
The $(\mathfrak{gl}_{n},\mathfrak{gl}_{m})$-duality, which we recall in Section 9, exchanges the usual Casimir connection with the Knizhnik-Zamolodchikov connection. Moreover, the monodromy of the KZ connection is equivalent to the quantum $R$-matrix by the Kohno–Drinfeld theorem [Drinfeld, Kohno]. Finally, by [ToledanoLaredoKohnoDrinfeldWeylGroup, Theorem 6.5], the action of the $R$-matrix for $U_{q}(\mathfrak{gl}_{n})$ is equal to the quantum Weyl group action for $U_{q}(\mathfrak{gl}_{m})$. These three observations are combined in the proof of [ToledanoLaredoKohnoDrinfeldWeylGroup, Theorem 7.1] to show that the monodromy of the usual Casimir connection is given by the quantum Weyl group action.
Under the orthogonal Howe duality of Section 10, the KZ connection for $\mathfrak{so}_{2n}$ is exchanged with the boundary Casimir connection for $\mathrm{O}_{m}\subset\mathrm{GL}_{m}$. The Kohno–Drinfeld theorem again implies that the monodromy of the KZ connection for $\mathfrak{so}_{2n}$ is equivalent to the $R$-matrix for $U_{q}(\mathfrak{so}_{2n})$. Finally, we computed, for $n\leq 4$, that under Howe duality the action of the $U_{q}(\mathfrak{so}_{2n})$ $R$-matrix is equal to the $U_{q}^{\iota}(\mathfrak{so}_{2})$ $\iota$quantum Weyl group operators [Zhang-Weinan-thesis].
@@ 9.1.1. Lie algebra and representations
We identify the Cartan subalgebra $\mathfrak{h}$ with $\mathbb{C}^{n}$ via the basis $\{E_{ii}|i=1\ldots n\}$ and denote by $\{\epsilon_{i}\}\subset\mathfrak{h}^{*}$ the dual basis. Then the roots are $\epsilon_{i}-\epsilon_{j}$ with the root spaces $\mathfrak{g}_{\epsilon_{i}-\epsilon_{j}}=\mathbb{C}\cdot E_{ij}$. The positive roots are $R_{+}=\{\epsilon_{i}-\epsilon_{j}|i<j\}$ with simple roots
@@
(9.1) $\mathrm{det}^{\alpha}:=\mathbb{C}\cdot\omega^{\alpha},\qquad E_{ij}\cdot\omega^{\alpha}=\delta_{ij}\alpha\cdot\omega^{\alpha}$
@@ 9.1.2. Skew polynomials
Denote by $\mathfrak{P}_{k}$ the space of anticommuting polynomials in $k$ variables $\{x_{i}|i=1\ldots k\}$. The algebra $\mathfrak{P}_{k}$ is $\mathbb{Z}_{\geq 0}$ graded, with generators $x_{i}$ in degree $1$. In what follows, we denote by $\mathbf{1}\in\mathfrak{P}_{n}$ the identity polynomial.
@@
(9.2) $E_{ij}\cdot g(x)=x_{i}\partial_{j}g(x),\ g(x)\in\mathfrak{P}_{k}$
@@ 9.3. Classical duality
$-E_{ij}E_{ji}+E_{ii}=-(\sum_{\alpha}x_{i\alpha}\partial_{j\alpha})(\sum_{\beta}x_{j\beta}\partial_{i\beta})+\sum_{\alpha}x_{i\alpha}\partial_{i\alpha}=-\sum_{\alpha,\beta}x_{i\alpha}\partial_{i\beta}x_{j\beta}\partial_{j\alpha}+\sum_{\alpha}x_{i\alpha}\partial_{i\alpha}.$
@@ 9.4. Quantum duality
In this subsection, we recall the quantum version of Section 9.3. For that, we will need to introduce the Yangian $\mathrm{Y}(\mathfrak{gl}_{m})$.
@@ 9.4.1. Yangian
which is an $\mathrm{End}(\mathbb{C}^{m}\otimes\mathbb{C}^{m})$-valued function of $u$, where $P=\sum_{\alpha,\beta}E_{\alpha\beta}\otimes E_{\beta\alpha}$ is the permutation operator. The Yangian $\mathrm{Y}(\mathfrak{gl}_{m})$ is generated by $t_{\alpha\beta}^{(k)}$ for $\alpha,\beta=1\ldots m$ and $k\geq 1$ subject to the relations
$[t_{\alpha\beta}^{(r+1)},t_{\gamma\delta}^{(s)}]-[t_{\alpha\beta}^{(r)},t_{\gamma\delta}^{(s+1)}]=t_{\gamma\beta}^{(r)}t_{\alpha\delta}^{(s)}-t_{\gamma\beta}^{(s)}t_{\alpha\delta}^{(r)}.$
@@
$t_{\alpha\beta}(u):=\delta_{\alpha\beta}+\sum_{k\geq 0}t_{\alpha\beta}^{(k)}u^{-k}\in\mathrm{Y}(\mathfrak{gl}_{m})[[u^{-1}]],$
@@
For any $\mathfrak{gl}_{m}$-representation $V$, we denote by $V(z)$ the pullback via the composition $\mathrm{ev}\circ\tau_{z}$. Observe that the automorphism $\tilde{\theta}$ of (9.8) lifts the Cartan involution $\theta$ of Section 5.
@@ 9.4.2. Duality
A quantum analog of the differential operators from Section 9.3 is as follows. Denote by $T_{z_{i}}$ the shift operator acting on meromorphic analytic functions $\mathbb{C}^{n}\rightarrow\mathfrak{P}_{nm}$ by
$[T_{z_{i}}f](z_{1},\ldots,z_{n})=f(z_{1},\ldots,z_{i}+\kappa,\ldots,z_{n}).$
@@ 10.1.1. Lie algebra $\mathfrak{so}_{k}$
Finite dimensional irreducible representations of $\mathfrak{so}_{2n}$ are described by sequences of half-integral numbers $\vec{l}=(l_{1},\ldots,l_{n})$ satisfying $l_{1}\geq\ldots\geq l_{n-1}\geq|l_{n}|$; we denote them by $V(\mathfrak{so}_{2n},\vec{l})$. The corresponding representation integrates to the group $\mathrm{SO}(2n)$ if and only if all numbers are integral, i.e. when $\mathbf{l}:=(l_{1},\ldots,l_{n-1},|l_{n}|)$ defines a partition.
@@
Finite dimensional irreducible representations of $\mathfrak{so}_{2n+1}$ are described by sequences $\vec{l}=(l_{1},\ldots,l_{n})$ of half-integral numbers such that $l_{1}\geq\ldots\geq l_{n}\geq 0$; we denote it by $V(\mathfrak{so}_{2n+1},\vec{l})$. This representation integrates to a representation of $\mathrm{SO}(2n+1)$ if and only if all the numbers are integral, i.e. $\mathbf{l}:=(l_{1},\ldots,l_{n})$ defines a partition.
@@ 10.1.2. Representations of $\mathrm{O}_{k}$
Recall the Cartan involution $\theta$ from Section 5. In the case of $\mathfrak{g}=\mathfrak{gl}_{k}$, it is given by $\theta(E_{\alpha\beta})=-E_{\beta\alpha}$ for all $\alpha,\beta\in\{1,\ldots,m\}$. In fact, it can be extended to an involution on the group $\mathrm{GL}_{k}$ of invertible matrices by $\theta(g)=(g^{-1})^{\mathrm{T}}$. Denote by $\mathrm{O}_{k}:=\mathrm{GL}^{\theta}_{k}$ the fixed point subgroup.
@@
To describe finite dimensional irreducible representations of $\mathrm{O}_{k}$, we use the root systems of Section 10.1.1. However, one must be careful when doing so. Note that $\mathrm{O}_{k}=\mathrm{GL}_{k}^{\theta}$ preserves the form on $\mathbb{C}^{k}$ such that $(e_{i},e_{j})=\delta_{ij}$. But the root systems in Section 10.1.1 correspond to Cartan subalgebras of $\mathfrak{so}_{k}$ which preserve a different bilinear form on $\mathbb{C}^{k}$ than the one preserved by $\mathfrak{gl}_{k}^{\theta}$.
When $k=2n+1$ is odd, then $-\mathrm{id}\in\mathrm{O}_{k}$ and $-\mathrm{id}\notin\mathrm{SO}_{k}$, so $\mathrm{O}_{k}$ is the direct product $\mathrm{SO}_{k}\times\mathbb{Z}/2$. In particular, all irreducible representations of $\mathrm{O}_{k}$ are isomorphic to the external tensor product of an irreducible representation of $\mathrm{SO}_{k}$ with an irreducible representation for $\mathbb{Z}/2$. For a partition $\mathbf{l}$, let $V(\mathrm{O}_{k},\mathbf{l})$ be the representation of $\mathrm{O}_{k}$ obtained from $V(\mathrm{SO}_{k},\mathbf{l})$ by making $-\mathrm{id}$ act trivially.
@@
Formally, denote by $V(\mathrm{O}_{k},\tilde{\mathbf{l}}):=V(\mathrm{SO}_{k},\mathbf{l})\otimes\det$, the representation of $\mathrm{O}_{k}$ obtained from $V(\mathrm{SO}_{k},\mathbf{l})$ by making $-\mathrm{id}$ act by $\mathrm{det}(-\mathrm{id})=-1$.
@@ 10.1.3. Spinor representation
$[\gamma^{\dagger}_{i},\gamma_{j}]_{+}=\delta_{ij},\qquad[\gamma^{\dagger}_{i},\gamma^{\dagger}_{j}]_{+}=[\gamma_{i},\gamma_{j}]_{+}=0,$
where $[-,-]_{+}$ is the anticommutator, i.e. $[x,y]_{+}=xy+yx$. The algebra $\mathrm{Cl}_{n}$ is $\mathbb{Z}/2$-graded by
@@
There is an action of $\mathrm{Cl}_{n}$ on skew polynomials $\mathfrak{P}_{n}$, see Section 9.1.2, defined by

### v3

@@ 8.1. Background on connections and differential operators
A flat connection is a conection $\nabla$ such that $\nabla_{\xi},\nabla_{\eta}]=\nabla_{[\xi,\eta]}$, for all vector fields $\xi,\eta$.
In coordinates, $\nabla$ is a flat connection if and only if $\nabla_{\partial/\partial z_{i}},\nabla_{\partial/\partial z_{j}}]=0$ for all $1\leq i,j\leq n$.
@@ 8.2. Boundary Casimir connection
In this subsection, we are in the setup of Section˜5. For instance, we consider an arbitrary simple Lie algebra $\mathfrak{g}$ with a Cartan involution $\theta$.
@@
One can easily see that Theorem˜8.7 follows from the following proposition.
@@
Let $N_{\alpha,\gamma}\in\mathbb{C}$ be the numbers such that $E_{\alpha},E_{\gamma}]=N_{\alpha,\gamma}E_{\alpha+\gamma}$. Then a summand of the expression above is equal to
@@
$T(\lambda,\mu),T(\nu,\mu)]=\sum_{\alpha,\beta\in R_{+}}\frac{\langle\alpha,\lambda\rangle\langle\beta,\nu\rangle}{\langle\alpha,\mu\rangle\langle\beta,\mu\rangle}[B^{2}_{\alpha},B^{2}_{\beta}]=\sum_{\begin{subarray}{c}\alpha,\beta\in R_{+}\\ \alpha\neq\beta\end{subarray}}\frac{\langle\alpha,\lambda\rangle\langle\beta,\nu\rangle}{\langle\alpha,\mu\rangle\langle\beta,\mu\rangle}[B^{2}_{\alpha},B^{2}_{\beta}].$
@@
The residue [^6] Recall for example that if $f$ is a rational function on $\mathbb{C}^{n}$, the residue along $z_{1}=0$ is computed as follows. Write $f=z_{1}^{-k}f_{-k}+z_{1}^{-k+1}f_{-k+1}+\dots$, where each $f_{i}$ is a rational function of $z_{2},\dots,z_{n}$, then $\mathrm{res}_{z_{1}=0}f=f_{-1}$. In particular, if $f$ has only a simple pole at $z_{1}=0$, i.e., if $f=z_{1}^{-1}f_{-1}+f_{0}+z_{1}f_{1}+\dots$, then $\mathrm{res}_{z_{1}=0}f=(z_{1}f)|_{z_{1}=0}$. at $\langle\alpha,\mu\rangle=0$ is given by
@@
by ˜8.9. Since the sum in equation (8.3) can presented as the sum of such terms, the residue is zero. We deduce that the function is regular.
It follows from Proposition˜3.5 and ˜3.15 that
@@
In particular, by Proposition˜8.4, we see that the boundary Casimir connection descends to the quotient $\mathfrak{h}^{\mathrm{reg}}/W$. Recall that the monodromy of the usual Casimir connection is given by the quantum Weyl group action [ToledanoLaredoQuantumWeyl]. Analogously, we propose the following conjecture.
@@
The $(\mathfrak{gl}_{n},\mathfrak{gl}_{m})$-duality, which we recall in Section˜9, exchanges the usual Casimir connection with the Knizhnik-Zamolodchikov connection. Moreover, the monodromy of the KZ connection is equivalent to the quantum $R$-matrix by the Kohno–Drinfeld theorem [Drinfeld, Kohno]. Finally, by [ToledanoLaredoKohnoDrinfeldWeylGroup, Theorem 6.5], the action of the $R$-matrix for $U_{q}(\mathfrak{gl}_{n})$ is equal to the quantum Weyl group action for $U_{q}(\mathfrak{gl}_{m})$. These three observations are combined in the proof of [ToledanoLaredoKohnoDrinfeldWeylGroup, Theorem 7.1] to show that the monodromy of the usual Casimir connection is given by the quantum Weyl group action.
[…]
@@ 9.1.1. Lie algebra and representations
We identify the Cartan subalgebra $\mathfrak{h}$ with $\mathbb{C}^{n}$ via the basis $\{E_{ii}|i=1,\ldots,n\}$ and denote by $\{\epsilon_{i}\}\subset\mathfrak{h}^{*}$ the dual basis. Then the roots are $\epsilon_{i}-\epsilon_{j}$ with the root spaces $\mathfrak{g}_{\epsilon_{i}-\epsilon_{j}}=\mathbb{C}\cdot E_{ij}$. The positive roots are $R_{+}=\{\epsilon_{i}-\epsilon_{j}|i<j\}$ with simple roots
@@
(9.1) $\mathrm{det}^{\alpha}:=\mathbb{C}\cdot\omega^{\alpha},\qquad E_{ij}\cdot\omega^{\alpha}=\delta_{ij}\alpha\cdot\omega^{\alpha}.$
@@ 9.1.2. Skew polynomials
Denote by $\mathfrak{P}_{k}$ the space of anticommuting polynomials in $k$ variables $\{x_{i}|i=1,\ldots,k\}$. The algebra $\mathfrak{P}_{k}$ is $\mathbb{Z}_{\geq 0}$ graded, with generators $x_{i}$ in degree $1$. In what follows, we denote by $\mathbf{1}\in\mathfrak{P}_{n}$ the identity polynomial.
@@
(9.2) $E_{ij}\cdot g(x)=x_{i}\partial_{j}g(x),\ g(x)\in\mathfrak{P}_{k}.$
@@ 9.3. Classical duality
$E_{ij}E_{ji}+E_{ii}=-(\sum_{\alpha}x_{i\alpha}\partial_{j\alpha})(\sum_{\beta}x_{j\beta}\partial_{i\beta})+\sum_{\alpha}x_{i\alpha}\partial_{i\alpha}=-\sum_{\alpha,\beta}x_{i\alpha}\partial_{i\beta}x_{j\beta}\partial_{j\alpha}+\sum_{\alpha}x_{i\alpha}\partial_{i\alpha}.$
@@ 9.4. Quantum duality
In this subsection, we recall the quantum version of Section˜9.3. For that, we will need to introduce the Yangian $\mathrm{Y}(\mathfrak{gl}_{m})$.
@@ 9.4.1. Yangian
which is an $\mathrm{End}(\mathbb{C}^{m}\otimes\mathbb{C}^{m})$-valued function of $u$, where $P=\sum_{\alpha,\beta}E_{\alpha\beta}\otimes E_{\beta\alpha}$ is the permutation operator. The Yangian $\mathrm{Y}(\mathfrak{gl}_{m})$ is generated by $t_{\alpha\beta}^{(k)}$ for $\alpha,\beta=1,\ldots,m$ and $k\geq 1$ subject to the relations
$t_{\alpha\beta}^{(r+1)},t_{\gamma\delta}^{(s)}]-[t_{\alpha\beta}^{(r)},t_{\gamma\delta}^{(s+1)}]=t_{\gamma\beta}^{(r)}t_{\alpha\delta}^{(s)}-t_{\gamma\beta}^{(s)}t_{\alpha\delta}^{(r)}.$
@@
$t_{\alpha\beta}(u):=\delta_{\alpha\beta}+\sum_{k>0}t_{\alpha\beta}^{(k)}u^{-k}\in\mathrm{Y}(\mathfrak{gl}_{m})[[u^{-1}]],$
@@
For any $\mathfrak{gl}_{m}$-representation $V$, we denote by $V(z)$ the pullback via the composition $\mathrm{ev}\circ\tau_{z}$. Observe that the automorphism $\tilde{\theta}$ of (9.8) lifts the Cartan involution $\theta$ of Section˜5.
@@ 9.4.2. Duality
A quantum analog of the differential operators from Section˜9.3 is as follows. Denote by $T_{z_{i}}$ the shift operator acting on meromorphic analytic functions $\mathbb{C}^{n}\rightarrow\mathfrak{P}_{nm}$ by
$T_{z_{i}}f](z_{1},\ldots,z_{n})=f(z_{1},\ldots,z_{i}+\kappa,\ldots,z_{n}).$
@@ 10.1.1. Lie algebra $\mathfrak{so}_{k}$
Finite dimensional irreducible representations of $\mathfrak{so}_{2n}$ are described by sequences of half-integral numbers $\vec{l}=(l_{1},\ldots,l_{n})$ satisfying $l_{1}\geq\ldots\geq l_{n-1}\geq|l_{n}|$; we denote them by $V(\mathfrak{so}_{2n},\vec{l})$. The corresponding representation integrates to the group $\mathrm{SO}(2n)$ if and only if all numbers are integral, i.e., when $\mathbf{l}:=(l_{1},\ldots,l_{n-1},|l_{n}|)$ defines a partition.
@@
Finite dimensional irreducible representations of $\mathfrak{so}_{2n+1}$ are described by sequences $\vec{l}=(l_{1},\ldots,l_{n})$ of half-integral numbers such that $l_{1}\geq\ldots\geq l_{n}\geq 0$; we denote it by $V(\mathfrak{so}_{2n+1},\vec{l})$. This representation integrates to a representation of $\mathrm{SO}(2n+1)$ if and only if all the numbers are integral, i.e., $\mathbf{l}:=(l_{1},\ldots,l_{n})$ defines a partition.
@@ 10.1.2. Representations of $\mathrm{O}_{k}$
Recall the Cartan involution $\theta$ from Section˜5. In the case of $\mathfrak{g}=\mathfrak{gl}_{k}$, it is given by $\theta(E_{\alpha\beta})=-E_{\beta\alpha}$ for all $\alpha,\beta\in\{1,\ldots,m\}$. In fact, it can be extended to an involution on the group $\mathrm{GL}_{k}$ of invertible matrices by $\theta(g)=(g^{-1})^{\mathrm{T}}$. Denote by $\mathrm{O}_{k}:=\mathrm{GL}^{\theta}_{k}$ the fixed point subgroup.
@@
To describe finite dimensional irreducible representations of $\mathrm{O}_{k}$, we use the root systems of Section˜10.1.1. However, one must be careful when doing so. Note that $\mathrm{O}_{k}=\mathrm{GL}_{k}^{\theta}$ preserves the form on $\mathbb{C}^{k}$ such that $(e_{i},e_{j})=\delta_{ij}$. But the root systems in Section˜10.1.1 correspond to Cartan subalgebras of $\mathfrak{so}_{k}$ which preserve a different bilinear form on $\mathbb{C}^{k}$ than the one preserved by $\mathfrak{gl}_{k}^{\theta}$.
When $k=2n+1$ is odd, then $\mathrm{id}\in\mathrm{O}_{k}$ and $\mathrm{id}\notin\mathrm{SO}_{k}$, so $\mathrm{O}_{k}$ is the direct product $\mathrm{SO}_{k}\times\mathbb{Z}/2$. In particular, all irreducible representations of $\mathrm{O}_{k}$ are isomorphic to the external tensor product of an irreducible representation of $\mathrm{SO}_{k}$ with an irreducible representation for $\mathbb{Z}/2$. For a partition $\mathbf{l}$, let $V(\mathrm{O}_{k},\mathbf{l})$ be the representation of $\mathrm{O}_{k}$ obtained from $V(\mathrm{SO}_{k},\mathbf{l})$ by making $\mathrm{id}$ act trivially.
@@
Formally, denote by $V(\mathrm{O}_{k},\tilde{\mathbf{l}}):=V(\mathrm{SO}_{k},\mathbf{l})\otimes\det$, the representation of $\mathrm{O}_{k}$ obtained from $V(\mathrm{SO}_{k},\mathbf{l})$ by making $\mathrm{id}$ act by $\mathrm{det}(-\mathrm{id})=-1$.
@@ 10.1.3. Spinor representation
$\gamma^{\dagger}_{i},\gamma_{j}]_{+}=\delta_{ij},\qquad[\gamma^{\dagger}_{i},\gamma^{\dagger}_{j}]_{+}=[\gamma_{i},\gamma_{j}]_{+}=0,$
where $[-,-]_{+}$ is the anticommutator, i.e., $x,y]_{+}=xy+yx$. The algebra $\mathrm{Cl}_{n}$ is $\mathbb{Z}/2$-graded by
@@
There is an action of $\mathrm{Cl}_{n}$ on skew polynomials $\mathfrak{P}_{n}$, see Section˜9.1.2, defined by